# EXCOGITO, an extensible coarse-graining toolbox for the investigation of biomolecules by means of low-resolution representations

### v1

@@ 4 Example application of EXCOGITO: icosalanine
Figure 4: (a-b): values of radius of gyration and root mean squared fluctuation (RMSF) extracted from the simulation. c: time evolution of mapping norms for $C_{\alpha}$ mapping and ${\bf M}^{nter}$ (see main text). d: comparison between the distribution of mapping entropy values of random (blue histogram) and optimized mappings (green histogram). e: time evolution of the cosine between ${\bf M}^{nter}$ and ${\bf M}^{cter}$ (see main text). f: probability of conserving each heavy atom in the icosolanine structure, displayed in the original, unrealistic, fully stretched conformation (1), and in three realistic conformations observed during the simulation (partially stretched (2), partially folded (3), and completely folded (4))
Given this trajectory, we first quantify (task measure) the mapping entropy for $500$ randomly extracted mappings with $40$ CG sites. We then follow the standard simulated annealing protocol (task optimize) to minimize $S_{map}$ over $48$ independent optimizations. Fig. 4d reports the non-overlapping distributions of values of $S_{map}$ arising for the two sets of mappings, and Fig. 4f shows how the probability of retaining each atom in the optimized solutions is unevenly distributed over the polypeptide chain. The two terminal regions are highly conserved by the minimization procedure, while the central region is more coarse-grained, especially in its $C_{\beta}$ atoms. Given the analysed set of configurations and values of energy, this suggests that the optimal CG mapping should assign higher resolution to the two terminal regions of the peptide, with a coarser description of the central region that only retains some backbone atoms.
@@
In Fig. 4c we aim at elucidating the behavior of the mapping norm (task norm, Eq. 22) for two markedly different mapping operators throughout this system’s trajectory. The first operator is the ${\bf M}^{C_{\alpha}}$ mapping, obtained by retaining only the $C_{\alpha}$ atoms of the system, while the second, ${\bf M}^{Nter}$, contains only the first $20$ atoms of the chain starting from the N terminal. Intuitively, the first mapping is very uniformly distributed over the protein, while the second is extremely localised in a specific region. From the plot we observe how $\mathcal{E}({\bf M}^{Nter})$ is consistently higher than $\mathcal{E}({\bf M}^{C_{\alpha}})$, as expected given its higher globularity. Moreover, $\mathcal{E}({\bf M}^{Nter})$ does not display relevant fluctuations, as the atoms retained by ${\bf M}^{Nter}$ do not change their mutual distances appreciably during the simulation. Instead, the ${\bf M}^{C_{\alpha}}$ mapping induces very wide fluctuations in $\mathcal{E}$, due to the continuous folding and unfolding of the polypeptide.
@@
Finally, Fig. 4e shows the cosine (task cosine, Eq. 19) of the angle between two mappings with $N=20$, namely ${\bf M}^{Nter}$ (see above) and ${\bf M}^{Cter}$, which contains only atoms coming from the C terminal region of the peptide. The cosine is very low when the polypeptide is in a stretched conformation and the mappings are therefore almost orthogonal. Instead, it approaches $1.0$ when the polypeptide is in a packed conformation (such as Fig. 4f (4)) and atoms of the two mappings are very close to each other, giving rise to an almost perfect parallelism.
@@ 5 Conclusions
Recent works 42, 31, 35, 3, 36 have emphasized the fundamental importance of the mapping between high- and low-resolution descriptions, whose origins are to be found in the renormalisation group approach to critical phenomena. While significant efforts have been directed towards developing CG force fields, one can observe in the literature a notable gap in the examination of the properties of mappings themselves, which nonetheless a few authors have started to address. The concept of mapping entropy, defined as the divergence between high-resolution and low-resolution configurations, has emerged as an important measure of the information retained by a mapping. The minimization of mapping entropy, in fact, offers valuable insights into the system’s function, unveiling nontrivial knowledge about its physical, chemical, and potentially biological properties. Furthermore, mappings can be leveraged even in the absence of sampled conformations, making use of ad hoc metrics to identify structural features and quantitatively explore the mapping space.
In this work we have presented EXCOGITO, a suite of routines that enables the analysis of macromolecular systems making use of the properties of mappings, most notably by means of quantities such as mapping entropy and mapping metrics. EXCOGITO is a toolbox software platform implemented in C and freely available from a public repository, and stands as a freely available, user-friendly tool empowering researchers to effectively explore the properties of complex biological or artificial macromolecules through the lens of low-resolution representations. By making this software available to the community, we hope to contribute to the field of soft and biological matter modelling, and facilitate further advancements in understanding complex molecular systems.
@@ 6.1 Launching EXCOGITO: mandatory files and external parameters
  • List of parameters of EXCOGITO. In the mandatory column, all (resp. no) indicates parameters that are always (resp. never) mandatory, while O, R, and M refer to parameters that are mandatory only for optimize, random, and measure (including the _kl counterparts) tasks, respectively. C0, C1, C2, C3, C4 correspond to the different clustering criteria (Sec. 6.2): for example, if the selected criterion is $2$, parameters min_nclust and max_nclust must be present.
@@ 6.2 Comparing coarse-grained structures and clustering the conformational space
$\text{RMSD}^{\text{CG}}({\bf M}({\bf x}),{\bf M}({\bf y}))=\sqrt{\frac{1}{N}\sum_{I=1}^{N}({\bf M}_{I}({\bf x})-\mathcal{RT}^{\text{CG}}{\bf M}_{I}({\bf y}))^{2}}$ (26)
@@
    as in Refs. 31, 44, with the purpose of increasing the robustness of the Simulated Annealing procedure employed in the mapping optimization. Here, $\mathcal{K}$ is the set of integers employed and ${S}_{map}^{K}$ is the mapping entropy associated with a specific choice of the number of clusters $K$.
@@
  • $\text{{criterion}}=3$ A fast version of $\text{{criterion}}=0$ that can be used only when a continuous trajectory is provided in input. In this case, the algorithm computes the pairwise $\text{RMSD}^{\text{CG}}$ matrix between a subset of the overall configurations of the trajectory, that is, one every stride (Tab. 2) configurations. For example, if $\text{{frames}}=101$ and $\text{{stride}}=50$, only “pivot” configurations number $1$, $51$ and $101$ are considered in the pairwise alignments. Subsequently, standard hierarchical clustering applied to this reduced matrix assigns the coarse-grained macrostate to each pivot configuration. Then, the remaining data points are labelled using a simple prescription: if the previous and following pivot configurations possess the same label, the latter is assigned to all the intermediate structures. Instead, if the two pivot points have been labelled differently by the clustering algorithm, each intermediate structure is assigned to the same cluster of the closest pivot, that is, the one corresponding to the lowest $\text{RMSD}^{\text{CG}}$. This approximation guarantees a substantial speed-up to the overall calculation, as the computation of the $\text{RMSD}^{\text{CG}}$ matrix and the following clustering are the most cumbersome tasks, scaling quadratically with the number of frames of the trajectory. More specifically, given a certain value of frames, $f$, and stride, $s$, the overall number of pairwise alignments, $N_{a}$, in the worst case scenario is given by:
@@
Figure 5: Schematic representation of criteria 0, 11, and 22 for conformational clustering. These are equivalent in the first stage of the procedure, where a $\text{RMSD}^{\text{CG}}$ matrix is calculated between all the configurations (frames, see Tab. 2) of a full-atom MD trajectory, observed through the glasses of a CG mapping. From this typically large matrix, the full dendrogram is constructed using the average linkage prescription. Then, conformational clusters can be selected in three manners, namely 0) cutting the dendrogram when nclust (equal to 33 in this case) leaves are present; 11) cutting the dendrogram when a certain value of cophenetic distance (on the ordinate) is reached, irrespectively of the number of leaves; 22) applying the procedure 0 for a set of 55 evenly spaced values of the number of clusters ($\{2,3,4,5,6\}$ in this case), determined by parameters min_nclust and max_nclust (22 and $6$ in this figure).

### v2

@@ 4 Example application of EXCOGITO: icosalanine
Figure 5: Results of the EXCOGITO analyses performed on icosalanine. Panels (a-b): values of the radius of gyration and root mean squared fluctuations (RMSF) extracted from an MD simulation of the system. Panel c: time evolution of the mapping norms for the $C_{\alpha}$ and ${\bf M}^{nter}$ mappings (see main text). Panel d: comparison between the distribution of mapping entropy values of random (blue histogram) and optimized mappings (green histogram). Panel e: time evolution of the cosine between ${\bf M}^{nter}$ and ${\bf M}^{cter}$ (see main text). Panel f: probability of conserving each (heavy) atom in the icosolanine structure, displayed in the original, unrealistic, fully stretched conformation (1), and in three realistic conformations observed during the simulation—partially stretched (2), partially folded (3), and mainly folded (4).
Fig. 5a-b report, respectively, the radius of gyration of the system along the trajectory and the root mean square fluctuation (RMSF) of its atoms. During the simulation, the molecule is observed to undergo several conformational transitions, going back and forth between what are rather elongated, partially folded, and completely folded arrangements. This variability is made evident by the time evolution of the protein’s radius of gyration displayed in Fig. 5a, whose value ranges from $6$ to $18$ Å.
Also the RMSF presented in Fig. 5b is indicative of the molecule’s flexibility, with its central regions and termini fluctuating importantly. Icosalanine displays a quite unstructured, polymer-like behavior, in that it broadly explores the conformational space without collapsing into a stable native structure. Evidence of this is the fact that, despite the sequence homogeneity of the molecule, its RMSF results appear to be asymmetric around the central monomers of the sequence: this is in part because the N and C termini are chemically different, and they hence entertain distinct interactions with the remainder of the molecule and the solvent. On the other hand, the $200$ ns duration of the trajectory, which generally allows a good sampling of the native state for a similarly-sized, globular protein, is probably too short for icosalanine to explore the conformational space exhaustively. However, this is not an issue in the present case, since the objective of this simulation is not to investigate the properties of icosalanine, but rather to illustrate the methods of EXCOGITO.
@@
To demonstrate the usage of EXCOGITO in practice, we start by elucidating the behaviour of the squared mapping norm (norm tool, Eq. 22) for two markedly different mapping operators throughout this system’s trajectory, see Fig. 5c. The first operator is the ${\bf M}^{C_{\alpha}}$ mapping, obtained by retaining only the $C_{\alpha}$ atoms of the system, while the second, ${\bf M}^{Nter}$, contains only the first $20$ atoms of the chain starting from the N terminal. Intuitively, the first mapping is very uniformly distributed over the protein, while the second is extremely localised in a specific region. From the plot we observe how $\mathcal{E}({\bf M}^{Nter})$ is consistently higher than $\mathcal{E}({\bf M}^{C_{\alpha}})$, as expected given its higher globularity. Moreover, $\mathcal{E}({\bf M}^{Nter})$ does not display relevant fluctuations, as the atoms retained by ${\bf M}^{Nter}$ do not change their mutual distances appreciably during the simulation. Instead, the ${\bf M}^{C_{\alpha}}$ mapping induces very wide fluctuations in $\mathcal{E}$, due to the continuous folding and unfolding of the polypeptide.
@@
As for the geometrical relation among different CG representations, Fig. 5e shows the cosine (cosine tool, Eq. 19) of the angle between two mappings with $N=20$, namely ${\bf M}^{Nter}$ (see above) and ${\bf M}^{Cter}$, which contains only atoms coming from the C terminal region of the peptide. The cosine is very low when the polypeptide is in a stretched conformation and the mappings are therefore almost orthogonal. Instead, it approaches $1.0$ when the polypeptide is in a packed conformation (such as Fig. 5f (4)) and atoms of the two mappings are very close to each other, giving rise to an almost perfect parallelism.
Finally, we quantify (measure tool) the mapping entropy for $500$ randomly extracted mappings with $40$ CG sites. We then follow the standard simulated annealing protocol (optimize tool) to minimize $S_{map}$ over $48$ independent optimizations. Fig. 5d reports the non-overlapping distributions of values of $S_{map}$ arising for the two sets of mappings, and Fig. 5f shows how the probability of retaining each atom in the optimized solutions is unevenly distributed over the polypeptide chain. The two terminal regions are highly conserved by the minimization procedure, while the central region is more coarse-grained, especially in its $C_{\beta}$ atoms. Given the analysed set of configurations and values of energy, this suggests that the optimal CG mapping should assign higher resolution to the two terminal regions of the peptide, with a coarser description of the central region that only retains some backbone atoms.
@@ 5 Conclusions
Recent works 42, 31, 35, 3, 36 have emphasized the fundamental importance of the mapping between high- and low-resolution descriptions of a system, whose origins are to be found in the renormalisation group approach to critical phenomena. While significant efforts have been directed towards developing CG force fields, one can observe in the literature a notable gap in the examination of the properties of mappings themselves, which nonetheless a few authors have started to address. The concept of mapping entropy, defined as the divergence between the reference, high-resolution configuration distribution and the one reconstructed from a low-resolution representation, has emerged as an important measure of the information retained by a mapping. The minimization of mapping entropy, in fact, offers valuable insights into the system’s function, unveiling nontrivial knowledge about its physical, chemical, and potentially biological properties. Furthermore, mappings can be leveraged even in the absence of sampled conformations, making use of ad hoc metrics to identify structural features and quantitatively explore the mapping space.
In this work we have presented EXCOGITO, a suite of routines that enables the analysis of macromolecular systems making use of the properties of mappings, most notably by means of quantities such as mapping entropy and mapping metrics. EXCOGITO is a toolbox software platform implemented in C and freely available from a public repository, and stands as a user-friendly tool empowering researchers to effectively explore the properties of complex biological or artificial macromolecules through the lens of low-resolution representations. By making this software available to the community, we hope to contribute to the field of soft and biological matter modelling, and facilitate further advancements in understanding complex molecular systems.
@@ 6.1 Launching EXCOGITO: mandatory files and external parameters
  • List of parameters of EXCOGITO. In the mandatory column, all (resp. no) indicates parameters that are always (resp. never) mandatory, while O, R, M, and D refer to parameters that are mandatory only for optimize, random, measure, and distance (including the _kl counterparts) tasks, respectively. C0, C1, C2, C3, C4 correspond to the different clustering criteria (Sec. 6.2): for example, if the selected criterion is $2$, parameters min_nclust and max_nclust must be present.
@@ 6.2 Comparing coarse-grained structures and clustering the conformational space
$\text{RMSD}^{\text{CG}}({\bf M}({\bf x}),{\bf M}({\bf y}))=\sqrt{\frac{1}{N}\sum_{I=1}^{N}({\bf M}_{I}({\bf x})-\mathcal{RT}{\bf M}_{I}({\bf y}))^{2}}$ (26)
@@
    as in Refs. 31 and 44, with the purpose of increasing the robustness of the SA procedure employed in the mapping optimization. Here, $\mathcal{K}$ is the set of integers employed and ${S}_{map}^{K}$ is the mapping entropy associated with a specific choice of the number of clusters $K$.
@@
  • $\text{{criterion}}=3$ A fast version of $\text{{criterion}}=0$ that can be used only when a continuous trajectory of the system is provided in input. In this case, the algorithm computes the pairwise $\text{RMSD}^{\text{CG}}$ matrix between a subset of the overall configurations of the trajectory, that is, one every stride (Tab. 3) configurations. For example, if $\text{{frames}}=101$ and $\text{{stride}}=50$, only “pivot” configurations number $1$, $51$ and $101$ are considered in the pairwise alignments. Subsequently, standard hierarchical clustering applied to this reduced matrix assigns the coarse-grained macrostate to each pivot configuration. Then, the remaining data points are labelled using a simple prescription: if the previous and following pivot configurations possess the same label, the latter is assigned to all the intermediate structures. Instead, if the two pivot points have been labelled differently by the clustering algorithm, each intermediate structure is assigned to the same cluster of the closest pivot, that is, the one corresponding to the lowest $\text{RMSD}^{\text{CG}}$. This approximation guarantees a substantial speed-up to the overall calculation, as the computation of the $\text{RMSD}^{\text{CG}}$ matrix and the following clustering are the most cumbersome tasks, scaling quadratically with the number of frames of the trajectory. More specifically, given a certain value of frames, $f$, and stride, $s$, the overall number of pairwise alignments, $N_{a}$, in the worst case scenario is given by:
@@
Figure 6: Schematic representation of criteria 0, 11, and 22 for conformational clustering. These are equivalent in the first stage of the procedure, where a $\text{RMSD}^{\text{CG}}$ matrix is calculated between all the configurations (frames, see Tab. 3) of a full-atom MD trajectory, observed through the glasses of a CG mapping. From this typically large matrix, the full dendrogram is constructed using the average linkage prescription. Then, conformational clusters can be selected in three manners, namely (0) cutting the dendrogram when nclust (equal to 33 in this case) leaves are present; 11) cutting the dendrogram when a certain value of cophenetic distance (on the ordinate) is reached, irrespectively of the number of leaves; 22) applying the procedure (0( for a set of 55 evenly spaced values of the number of clusters ($\{2,3,4,5,6\}$ in this case), determined by parameters min_nclust and max_nclust (22 and $6$ in this figure).